Solving Systems by Graphing and Substitution
x+3y=2
-x+y=1

Answer by ewatrrr(24785)


Hi,
x+3y=2 OR y = -(1/3)x + 2/3
-x+y=1 OR y = x + 1
4y = 3
y = 3/4
-x + 3/4= 1
3/4-1 = x
-1/4 = x
Lines intersect at Pt(-1/4,3/4)